Your parents are buying a house for $187,500. They have a good credit rating, are making a 20% down payment, and expect to pay $1,575/month. The interest rate for the mortgage is 4.65%. How much interest is paid at the end of the second month?


1
Expert's answer
2021-05-27T13:08:17-0400

The first instalment:

"187500\cdot20\%=\$\ 37500"

The loan amount will be:

"187500-37500=\$\ 150000"

Monthly rate will be:

"4.65\%/12=0.3875\%"

The interest rate for the first month is: 

"150000\cdot0.3875\%=\$\ 581.25"

The loan repayment amount is:

"1575-581.25=\$\ 993.75"

At the end of the month, parents owe: 

"150000-993.75=\$\ 149006.25"

During the second month the amount of payment, which goes to interest, is:

"149006.25\cdot0.3875=\$\ 577.40"


Interest which is paid at the end of the second month:

"1575-577.40=\$\ 997.60"